OpenStudy (paul10175):

ex − 12e−x − 1 = 0 solve to 4 decimal places. please

OpenStudy (paul10175):

the x values are exponents

OpenStudy (anonymous):

this? \[e^x-12e^{-x}-1=0\]

OpenStudy (paul10175):

yes peachpi

OpenStudy (anonymous):

Multiply through by \(e^x\). \(e^{2x}-12-e^x=0\) rearrange and factor \(e^{2x}-e^x-12=0\) \((e^x-4)(e^x+3)=0\)

OpenStudy (paul10175):

hell yea that was right. I just solved for x, but e^x+3 is not in the domain